# Blind Baselines Beat Membership Inference Attacks for Foundation Models

### v1

@@ 4.2.3 Project Gutenberg
This dataset Meeus et al. (2023) consists of books from Project Gutenberg. The members are books contained in the RedPajama dataset Together (2023), more specifically the set of “PG-19” books collected by Rae et al. (2019) which were all originally published before 1919. The non-members consist of books added to the Project Gutenberg repository after the creation of the PG-19 corpus in February 2019. To mitigate the obvious distribution shift in the publication years of members and non-members, Meeus et al. (2023) filter both the members and non-members to only contain books published between 1850 and 1910 (i.e., members are books published between 1850 and 1910, which were added to Project Gutenberg before February 2019. Non-members are books that were originally published in the same period, but added to Project Gutenberg after February 2019). The authors evaluate a new MI attack (see Section 4.1.4 for details) against the OpenLLaMA model Geng & Liu (2023) trained on RedPajama.
@@
While the publication dates of the members and non-members are similarly distributed, we find that the date on which a book is added to Project Gutenberg still introduces a noticeable distribution shift. Indeed, it appears that the format of the preface metadata that Project Gutenberg adds to books has changed at some point between the collection of PG-19 and the non-member corpus (e.g., the link to the HTML version of a book used to have a .htm file extension, and this was changed to a .html extension for books added after a certain date). By exploiting such small discrepancies, our greedy n-gram attack applied to the first 1,000 characters of a book achieves 59.6% TPR at a 1% FPR, 3 times higher than the best MI attack evaluated on this dataset.
Of course, it may be possible to preprocess this particular dataset to get rid of these formatting discrepancies. Nevertheless, we believe our results illustrate how brittle the process of re-collecting an identically distributed non-member set can be. Moreover, even if we ignore any formatting information, and apply our attack directly to the book text, we can still distinguish members and non-members far better than chance, with a TPR of 16.6% at a 1% FPR.
@@ 5 A Path Forward: the Pile, DataComp and DataComp-LM
An alternative avenue is to forego MI evaluations on arbitrary foundation models, and focus on those models for which identically distributed sets of members and non-members are available. A prime example is the Pile Gao et al. (2020), which has an official test set. Models trained on the Pile training set—e.g., Pythia Biderman et al. (2023) or GPT-NeoX-20B citeblack2022gpt—are thus a popular target for MI attack evaluations (see e.g., Duan et al. (2024); Maini et al. (2024); Li et al. (2023)). Unfortunately, the Pile only contains text, and is too small to train state-of-the-art language models.
@@
Table 4: Models trained on datasets with random train-test splits for DataComp Gadre et al. (2024) and DataComp-LM Li et al. (2024). For each dataset pool, we report the pool’s absolute size (in image-text pairs, respectively tokens) and its relative size compared to the global pool that it was randomly sampled from. Gadre et al. (2024) and Li et al. (2024) have pretrained ViTs and LLMs on each pool size.
@@
Future web-scale dataset and model developers could similarly release an official IID split to enable rigorous MI evaluations for a wider variety of models and modalities.

### v2

@@ 4.2.2 Project Gutenberg
This dataset Meeus et al. (2023) consists of books from Project Gutenberg. The members are books contained in the RedPajama dataset Together (2023), more specifically the set of ‘‘PG-19’’ books collected by Rae et al. (2019) which were all originally published before 1919. The non-members consist of books added to the Project Gutenberg repository after the creation of the PG-19 corpus in February 2019. To mitigate the obvious distribution shift in the publication years of members and non-members, Meeus et al. (2023) filter both the members and non-members to only contain books published between 1850 and 1910 (i.e., members are books published between 1850 and 1910, which were added to Project Gutenberg before February 2019. Non-members are books that were originally published in the same period, but added to Project Gutenberg after February 2019). The authors evaluate a new MI attack (see Section 4.1.4 for details) against the OpenLLaMA model Geng & Liu (2023) trained on RedPajama.
@@
While the publication dates of the members and non-members are similarly distributed, we find that the date on which a book is added to Project Gutenberg still introduces a noticeable distribution shift. Indeed, it appears that the format of the preface metadata that Project Gutenberg adds to books has changed at some point between the collection of PG-19 and the non-member corpus (e.g., the link to the HTML version of a book used to have a .htm file extension, and this was changed to a .html extension for books added after a certain date). By exploiting such small discrepancies, our greedy n-gram attack applied to the first 1,000 characters of a book achieves $55.1\%$ TPR at a $1\%$ FPR, 3 times higher than the best MI attack evaluated on this dataset.
Of course, it may be possible to preprocess this particular dataset to get rid of these formatting discrepancies. Nevertheless, we believe our results illustrate how brittle the process of re-collecting an identically distributed non-member set can be. Moreover, even if we ignore any formatting information, and apply our attack directly to the book text, we can still distinguish members and non-members far better than chance, with a TPR of $16.6\%$ at a $1\%$ FPR.
@@ 5 A Path Forward: the Pile, DataComp and DataComp-LM
An alternative avenue is to forego MI evaluations on arbitrary foundation models, and focus on those models for which identically distributed sets of members and non-members are available. A prime example is the Pile Gao et al. (2020), which has an official test set. Models trained on the Pile training set—e.g., Pythia Biderman et al. (2023) or GPT-NeoX-20B Black et al. (2022)—are thus a popular target for MI attack evaluations (see e.g., Duan et al. (2024); Maini et al. (2024); Li et al. (2023)). Unfortunately, the Pile only contains text, and is too small to train state-of-the-art language models.
@@
Таблица 4: Models trained on datasets with random train-test splits for DataComp Gadre et al. (2024) and DataComp-LM Li et al. (2024). For each dataset pool, we report the pool’s absolute size (in image-text pairs, respectively tokens) and its relative size compared to the global pool that it was randomly sampled from. Gadre et al. (2024) and Li et al. (2024) have pretrained ViTs and LLMs on each pool size.
@@
Future web-scale dataset and model developers could similarly release an official IID (independent and identically distributed) split to enable rigorous MI evaluations for a wider variety of models and modalities.
@@ 6 Conclusion
7 Reproducibility
Our results can be reproduced using our code-base available at: https://github.com/ethz-spylab/Blind-MIA. The code-base includes most of the datasets and scripts to download the datasets that were too large to include.
Список литературы
  • Biderman et al. (2023) Stella Biderman, Hailey Schoelkopf, Quentin Gregory Anthony, Herbie Bradley, Kyle O’Brien, Eric Hallahan, Mohammad Aflah Khan, Shivanshu Purohit, USVSN Sai Prashanth, Edward Raff, et al. Pythia: A suite for analyzing large language models across training and scaling. In International Conference on Machine Learning, pp.  2397–2430. PMLR, 2023.
  • Black et al. (2022) Sid Black, Stella Biderman, Eric Hallahan, Quentin Anthony, Leo Gao, Laurence Golding, Horace He, Connor Leahy, Kyle McDonell, Jason Phang, et al. Gpt-neox-20b: An open-source autoregressive language model. arXiv preprint arXiv:2204.06745, 2022.
  • Bommasani et al. (2021) Rishi Bommasani, Drew A Hudson, Ehsan Adeli, Russ Altman, Simran Arora, Sydney von Arx, Michael S Bernstein, Jeannette Bohg, Antoine Bosselut, Emma Brunskill, et al. On the opportunities and risks of foundation models. arXiv preprint arXiv:2108.07258, 2021.
  • Bourtoule et al. (2021) Lucas Bourtoule, Varun Chandrasekaran, Christopher A Choquette-Choo, Hengrui Jia, Adelin Travers, Baiwu Zhang, David Lie, and Nicolas Papernot. Machine unlearning. In 2021 IEEE Symposium on Security and Privacy (SP), pp.  141–159. IEEE, 2021.
  • Carlini et al. (2021) Nicholas Carlini, Florian Tramer, Eric Wallace, Matthew Jagielski, Ariel Herbert-Voss, Katherine Lee, Adam Roberts, Tom Brown, Dawn Song, Ulfar Erlingsson, et al. Extracting training data from large language models. In 30th USENIX Security Symposium (USENIX Security 21), pp.  2633–2650, 2021.
  • Carlini et al. (2022) Nicholas Carlini, Steve Chien, Milad Nasr, Shuang Song, Andreas Terzis, and Florian Tramer. Membership inference attacks from first principles. In 2022 IEEE Symposium on Security and Privacy (SP), pp.  1897–1914. IEEE, 2022.
  • Duan et al. (2024) Michael Duan, Anshuman Suri, Niloofar Mireshghallah, Sewon Min, Weijia Shi, Luke Zettlemoyer, Yulia Tsvetkov, Yejin Choi, David Evans, and Hannaneh Hajishirzi. Do membership inference attacks work on large language models? arXiv preprint arXiv:2402.07841, 2024.
  • Duarte et al. (2024) André V Duarte, Xuandong Zhao, Arlindo L Oliveira, and Lei Li. De-cop: Detecting copyrighted content in language models training data. arXiv preprint arXiv:2402.09910, 2024.
  • Dubiński et al. (2024) Jan Dubiński, Antoni Kowalczuk, Stanisław Pawlak, Przemyslaw Rokita, Tomasz Trzciński, and Paweł Morawiecki. Towards more realistic membership inference attacks on large diffusion models. In Proceedings of the IEEE/CVF Winter Conference on Applications of Computer Vision, pp.  4860–4869, 2024.
  • Engstrom et al. (2020) Logan Engstrom, Andrew Ilyas, Shibani Santurkar, Dimitris Tsipras, Jacob Steinhardt, and Aleksander Madry. Identifying statistical bias in dataset replication. In International Conference on Machine Learning, pp.  2922–2932. PMLR, 2020.
  • Gadre et al. (2024) Samir Yitzhak Gadre, Gabriel Ilharco, Alex Fang, Jonathan Hayase, Georgios Smyrnis, Thao Nguyen, Ryan Marten, Mitchell Wortsman, Dhruba Ghosh, Jieyu Zhang, et al. DataComp: In search of the next generation of multimodal datasets. Advances in Neural Information Processing Systems, 36, 2024.
  • Gao et al. (2020) Leo Gao, Stella Biderman, Sid Black, Laurence Golding, Travis Hoppe, Charles Foster, Jason Phang, Horace He, Anish Thite, Noa Nabeshima, et al. The Pile: An 800GB dataset of diverse text for language modeling. arXiv preprint arXiv:2101.00027, 2020.
  • Gemini Team (2023) Gemini Team. Gemini: a family of highly capable multimodal models. arXiv preprint arXiv:2312.11805, 2023.
  • Geng & Liu (2023) Xinyang Geng and Hao Liu. OpenLLaMA: An open reproduction of LLaMA, May 2023. URL https://github.com/openlm-research/open_llama.
  • Jayaraman et al. (2020) Bargav Jayaraman, Lingxiao Wang, Katherine Knipmeyer, Quanquan Gu, and David Evans. Revisiting membership inference under realistic assumptions. arXiv preprint arXiv:2005.10881, 2020.
  • Ko et al. (2023) Myeongseob Ko, Ming Jin, Chenguang Wang, and Ruoxi Jia. Practical membership inference attacks against large-scale multi-modal models: A pilot study. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pp.  4871–4881, 2023.
  • Li et al. (2024) Jeffrey Li, Alex Fang, Georgios Smyrnis, Maor Ivgi, Matt Jordan, Samir Gadre, Hritik Bansal, Etash Guha, Sedrick Keh, Kushal Arora, et al. DataComp-LM: In search of the next generation of training sets for language models. arXiv preprint arXiv:2406.11794, 2024.
  • Li et al. (2023) Marvin Li, Jason Wang, Jeffrey Wang, and Seth Neel. MoPe: Model perturbation-based privacy attacks on language models. arXiv preprint arXiv:2310.14369, 2023.
  • Maini et al. (2024) Pratyush Maini, Hengrui Jia, Nicolas Papernot, and Adam Dziedzic. LLM dataset inference: Did you train on my dataset?, 2024.
  • Meeus et al. (2023) Matthieu Meeus, Shubham Jain, Marek Rei, and Yves-Alexandre de Montjoye. Did the neurons read your book? Document-level membership inference for large language models. arXiv preprint arXiv:2310.15007, 2023.
  • Meeus et al. (2024a) Matthieu Meeus, Shubham Jain, Marek Rei, and Yves-Alexandre de Montjoye. Inherent challenges of post-hoc membership inference for large language models. arXiv preprint arXiv:2406.17975, 2024a.
  • Meeus et al. (2024b) Matthieu Meeus, Igor Shilov, Manuel Faysse, and Yves-Alexandre de Montjoye. Copyright traps for large language models. arXiv preprint arXiv:2402.09363, 2024b.
  • OpenAI (2023) OpenAI. GPT-4 technical report. arXiv preprint arXiv:2303.08774, 2023.
  • Oren et al. (2023) Yonatan Oren, Nicole Meister, Niladri Chatterji, Faisal Ladhak, and Tatsunori B Hashimoto. Proving test set contamination in black box language models. arXiv preprint arXiv:2310.17623, 2023.
  • Panaitescu-Liess et al. (2024) Michael-Andrei Panaitescu-Liess, Zora Che, Bang An, Yuancheng Xu, Pankayaraj Pathmanathan, Souradip Chakraborty, Sicheng Zhu, Tom Goldstein, and Furong Huang. Can watermarking large language models prevent copyrighted text generation and hide training data? arXiv preprint arXiv:2407.17417, 2024.
  • Radford et al. (2019) Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, Ilya Sutskever, et al. Language models are unsupervised multitask learners. OpenAI blog, 1(8):9, 2019.
  • Rae et al. (2019) Jack W Rae, Anna Potapenko, Siddhant M Jayakumar, and Timothy P Lillicrap. Compressive transformers for long-range sequence modelling. arXiv preprint arXiv:1911.05507, 2019.
  • Raffel et al. (2020) Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter J Liu. Exploring the limits of transfer learning with a unified text-to-text transformer. Journal of machine learning research, 21(140):1–67, 2020.
  • Ramesh et al. (2021) Aditya Ramesh, Mikhail Pavlov, Gabriel Goh, Scott Gray, Chelsea Voss, Alec Radford, Mark Chen, and Ilya Sutskever. Zero-shot text-to-image generation. In International Conference on Machine Learning, 2021.
  • Recht et al. (2019) Benjamin Recht, Rebecca Roelofs, Ludwig Schmidt, and Vaishaal Shankar. Do ImageNet classifiers generalize to ImageNet? In International conference on machine learning, pp.  5389–5400. PMLR, 2019.
  • Schuhmann et al. (2022) Christoph Schuhmann, Romain Beaumont, Richard Vencu, Cade Gordon, Ross Wightman, Mehdi Cherti, Theo Coombes, Aarush Katta, Clayton Mullis, Mitchell Wortsman, et al. LAION-5B: An open large-scale dataset for training next generation image-text models. Advances in Neural Information Processing Systems, 35:25278–25294, 2022.
  • Shi et al. (2023) Weijia Shi, Anirudh Ajith, Mengzhou Xia, Yangsibo Huang, Daogao Liu, Terra Blevins, Danqi Chen, and Luke Zettlemoyer. Detecting pretraining data from large language models. arXiv preprint arXiv:2310.16789, 2023.
  • Shokri et al. (2017) Reza Shokri, Marco Stronati, Congzheng Song, and Vitaly Shmatikov. Membership inference attacks against machine learning models. In 2017 IEEE symposium on security and privacy (SP), pp.  3–18. IEEE, 2017.
  • Soldaini et al. (2024) Luca Soldaini, Rodney Kinney, Akshita Bhagia, Dustin Schwenk, David Atkinson, Russell Authur, Ben Bogin, Khyathi Chandu, Jennifer Dumas, Yanai Elazar, et al. Dolma: An open corpus of three trillion tokens for language model pretraining research. arXiv preprint arXiv:2402.00159, 2024.
  • Together (2023) Together. RedPajama: an open dataset for training large language models, 2023. URL https://github.com/togethercomputer/RedPajama-Data.
  • Touvron et al. (2023) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. LLaMA: Open and efficient foundation language models. arXiv preprint arXiv:2302.13971, 2023.
  • Yeom et al. (2018) Samuel Yeom, Irene Giacomelli, Matt Fredrikson, and Somesh Jha. Privacy risk in machine learning: Analyzing the connection to overfitting. In 2018 IEEE 31st computer security foundations symposium (CSF), pp.  268–282. IEEE, 2018.
  • Zhang et al. (2024a) Jingyang Zhang, Jingwei Sun, Eric Yeats, Yang Ouyang, Martin Kuo, Jianyi Zhang, Hao Yang, and Hai Li. Min-K%++: Improved baseline for detecting pre-training data from large language models. arXiv preprint arXiv:2404.02936, 2024a.
  • Zhang et al. (2024b) Weichao Zhang, Ruqing Zhang, Jiafeng Guo, Maarten de Rijke, Yixing Fan, and Xueqi Cheng. Pretraining data detection for large language models: A divergence-based calibration method. arXiv preprint arXiv:2409.14781, 2024b.
Приложение A Appendix
Рис. 1: In some cases, the distribution shift is easy to visualise. Here is the PCA plot of the WikiMIA dataset which shows members and non-members forming distinguishable clusters.